Prove that n^4+4^n is not a prime number for any integer n greater than 1
1
Expert's answer
2020-01-03T11:51:49-0500

Consider 4n+n4;

"\forall"n, n"\in" N /{1} it is true that

n=2m+1 or n=2m.

If n=2m+1,

then 4(2m+1)+(2m+1)4 = 22(2m+1)+ ((2m+1)2)2 = 22(2m+1)+ ((2m+1)2)2 + 2 ×22m+1 ×(2m+1)2 - 2 ×22m+1 ×(2m+1)2= (22m+1 + (2m+1)2)2 - 22m+2 × (2m+1)² = (22m+1 + (2m+1)2)2 - (2m+1×(2m+1))2=(22m+1+(2m+1) - 2m+1×(2m+1))(22m+1+(2m+1)+2m+1×(2m+1)) is a product of two multipliers, they are not 1 simultaneously, hence the product is not prime.

If n=2m,

then

42m+(2m)4 is an even number because

42m is an even number "\forall" m, m"\in" N,

(2m)4 is an even number "\forall" m, m"\isin" N.

Thus, 42m+(2m)4 is an even number.